Please help I am not understanding:
-6(x+1)+4=16
(-6x+6)+4-4=16-4=12
x=12
??
Thank you in advance

Answer by stanbon(75887)

-6(x+1)+4=16
-6x-6+4 = 16
-6x-2 = 16
-6x = 18
x = -3
===============
Cheers,
Stan H.